OpenStudy (anonymous):

HELP MEH! PLS! Convert x-4y=-145.28 to point-slope. Ill give medals!

OpenStudy (triciaal):

to slope-intercept y = mx + b when x = 0 what is y? when y = 0 what is x? you have the slope, m for any point (x, y) point-slope formula y-y1= m(x-x1) where (x1,y1) is a point on that line .